OpenStudy (anonymous):

sorry i cant seem to get it im dumb What is the solution to this equation? 2x - 5 + 3x = 35 A. x = 20 B. x = 15 C. x = 6 D. x = 8

OpenStudy (anonymous):

2x - 5 + 3x = 35 5x=35+5=40 x=40/5=8 so D is correct

OpenStudy (anonymous):

yes d

OpenStudy (anonymous):

i be do da problem uuumm

OpenStudy (anonymous):

2x-5+3x = 35, retrice5 to each side. and add 2+3 x, what does the problem look like now?

OpenStudy (anonymous):

Simplifying 2x + -5 + 3x = 35 Reorder the terms: -5 + 2x + 3x = 35 Combine like terms: 2x + 3x = 5x -5 + 5x = 35 Solving -5 + 5x = 35 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5' to each side of the equation. -5 + 5 + 5x = 35 + 5 Combine like terms: -5 + 5 = 0 0 + 5x = 35 + 5 5x = 35 + 5 Combine like terms: 35 + 5 = 40 5x = 40 Divide each side by '5'. x = 8 Simplifying x = 8